# Selecting informative conformal prediction sets with false coverage rate control

### v1

@@ 3.2 Informative selective conditional prediction sets (InfoSCOP)
Throughout this section, we consider the iid model. It turns out that InfoSP can be too conservative in some contexts; this is manifest in the true FCR being much smaller than the nominal $\alpha$ level, see illustrations in § 4 and § 5. To avoid this, we can adapt the conditional approach of Bao et al., (2024) to our framework and start by selecting test samples and calibration samples. We would like, following the initial selection rule to have as few as possible test samples for which $\mathcal{I}$-informative prediction sets cannot be constructed. Specifically, we follow the following steps
@@
    For $i+n\in\mathcal{S}^{(0)}\cap[n+1,n+m]$, compute the conformal $p$-values using calibration set $\{(X_{j},Y_{j}),j\in\mathcal{S}^{(0)}\cap[r+1,n]\}$:
@@
Also, we assume that the initial selection $\mathcal{S}^{(0)}$ satisfies the following permutation preserving assumption.
@@
The initial selection $\mathcal{S}^{(0)}$ is typically the result of a multiple testing procedure that is applied to the examples in $[r+1,n+m]$, that uses $p$-values computed using $((X_{j},Y_{j}),j\in[r])$ as calibration sample and $(X_{i},i\in[r+1,n+m])$ as covariate test sample, which immediately satisfies Assumption 3. We can use $\mathrm{BH}(\mathbf{q})$ as an initial selection stage (where the $q_{i}$’s are computed with the aforementioned sample split) so that selected examples are likely to correspond to $Y_{n+i}$ where an informative prediction set can be built. For excluding a null range in regression or a null class in classification, another choice is to use an appropriate BH procedure for testing that the examples from $[r+1,n+m]$ are from that null.
@@
The informative selective conditional prediction set procedure pre-processed with the initial selection rule $\mathcal{S}^{(0)}$, called InfoSCOP, is defined as the InfoSP procedure of Definition 3 applied with the pre-processed $p$-value family $\mathbf{\overline{p}}^{0}$ (24), that is,
$\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSCOP}}}(\mathbf{\overline{p}})=\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{\overline{p}}^{0})=(\mathcal{C}^{\alpha^{0}}_{n+i}(\mathbf{\overline{p}}^{0}))_{i\in\mathrm{BH}(\mathbf{q}^{0})},$
where $\alpha^{0}=\alpha|\mathrm{BH}(\mathbf{q}^{0})|/|\mathcal{S}^{(0)}\cap[n+1,n+m]|$ and the $\mathcal{I}$-adjusted $p$-values $\mathbf{q}^{0}$ are computed via (11) from the pre-processed $p$-values $\mathbf{\overline{p}}^{0}$.
@@
The proof is provided in § C.2.
As we will see in the next sections, while it maintains the FCR guarantee, InfoSCOP can greatly improve over InfoSP. The main reason is that adjusting for selection is cheaper after the initial selection step: by reducing the fraction of examples in the test sample for which it is not possible to construct $\mathcal{I}$ informative prediction sets, the correction term $\alpha^{0}$ is expected to be close to $\alpha$ (or not much smaller than $\alpha$). Another reason is that the selection-conditional $p$-values $\mathbf{\overline{p}}^{0}$ will be better in settings where the initial selection step tends to remove the examples from the calibration set that have large non-conformity scores, see § 4, § E and § F for such cases.
In general, the way InfoSCOP can improve over InfoSP depends on the context. For instance, for excluding a null range in classification, a null class in regression or for length restriction in regression, we show in § 4 and § 5, respectively, the great potential advantage of using the initial selection. On the other hand, we also show that for non-trivial classification, there may be no advantage of initial selection (in fact, there can be a light disadvantage since the calibration sample after initial selection is smaller, as demonstrated in § 5.4).
@@ 4 Application to regression
This section is devoted to the regression case (that is, $\mathcal{Y}=\mathbb{R}$), as already introduced in Example 1.1. Throughout the section, we work in the iid model. Illustrations for other informative selections can be found in § E.
4.1 Excluding an interval $[a,b]$ of the prediction interval
We focus here on the case where the user wants to build prediction intervals only for labels such that $Y_{n+i}<a$ or $Y_{n+i}>b$, which corresponds to excluding the interval $\mathcal{Y}_{0}=[a,b]$ of the prediction interval, where $a<b$ are two benchmark values. This corresponds to common practice where users are interested only in reporting prediction intervals for individuals with “abnormal” outcomes. Setting $a=-\infty$ recover the case where we only want to report prediction interval for examples such that $Y_{n+i}>b$, which is the selection considered in Jin and Candes, (2023). We focus on two-sided prediction intervals here (the case of one-sided prediction intervals is postponed to § E). The choice of score function defines InfoSP and entails all the desired inferential guarantees. We formalize this for the locally weighted score function in Corollary 4.1. Using other score functions is also possible, e.g., the score function that corresponds to conformal quantile regression, see Remark 4.1.
@@ 4.1 Excluding an interval $[a,b]$ of the prediction interval
Consider the iid model in the regression case, consider the locally weighted score function $S_{y}(x)=|\mu(x)-y|/\sigma(x)$ and suppose that Assumption 1 holds. Then the following holds for InfoSP with informative collection =$\mathcal{I}=\{\mbox{$I$ interval of }\mathbb{R}\>:\>I\cap[a,b]=\emptyset\}$ and full calibrated $p$-value collection $\mathbf{\overline{p}}$ (7):
@@
    $q_{i}=\overline{p}^{(a)}_{i}{\mathbf{1}\{\mu(X_{n+i})<a\}}+\overline{p}^{(b)}_{i}{\mathbf{1}\{\mu(X_{n+i})>b\}}+{\mathbf{1}\{a\leq\mu(X_{n+i})\leq b\}},$ (25)
    where $\overline{p}^{(a)}_{i}$ and $\overline{p}^{(b)}_{i}$ are the full-calibrated $p$-values (7).
@@
    The selected prediction intervals do not contain $[a,b]$ and are of the form $\mathcal{C}_{n+i}=[\mu(x)-S_{(n_{\alpha}(\mathbf{p}))}\sigma(x),\mu(x)+S_{(n_{\alpha}(\mathbf{p}))}\sigma(x)]$, where $S_{(1)}\leq\dots\leq S_{(n)}$ are the ordered calibration scores $S_{Y_{j}}(X_{j})$, $1\leq j\leq n$ (with $S_{(n+1)}=+\infty$), and $n_{\alpha}(\mathbf{p})=\lceil(1-\alpha|\mathcal{S}(\mathbf{p})|/m)(n+1)\rceil$.
@@
Point (i) comes from Example 2.1 because $\min_{y\in[a,b]}S_{y}(x)$ is 00 if $\mu(x)\in[a,b]$, $S_{a}(x)$ if $\mu(x)<a$, and $S_{b}(x)$ if $\mu(x)>b$. Point (ii) comes from Lemma 2.1, (iii) from Remark 2.1 and the fact that $q_{i}\leq\alpha|\mathcal{S}(\mathbf{p})|/m$ iff $\mathcal{C}_{n+i}$ does not contain $[a,b]$. Point (iv) comes from Theorem 3.1. ∎
Hence, our method, in addition to providing an FCR control on the selected (iv) ensures that the obtained prediction intervals are informative in the sense that they do not include benchmark values (i.e., values in $[a,b]$). This ensures that the selection method is meaningful for the considered aim, which formally entails the FDR control (ii). Obviously, since InfoSCOP is an InfoSP method for preprocessed $p$-values (24), a similar result holds for InfoSCOP, for any initial selection step $\mathcal{S}^{(0)}\subset[r+1,n+m]$ that satisfies the permutation preserving Assumption 3.
Corollary 4.1 is illustrated on Figure 3 when $\mathcal{S}^{(0)}$ is taken here as $\mathrm{BH}(\mathbf{q})$ at level $2\alpha$ (with the score $S_{y}(x)=|\mu(x)-y|/\sigma(x)$). In the first row, errors appear more often near to the selection area. Hence, while the marginal prediction intervals control the FCR at level $\alpha$ when selecting all the covariates (as granted by (9)), the FCR is inflated for a naive selection that selects each example with a prediction interval at level $\alpha$ not intersecting $[a,b]$ (that is, naive selection is given by $\mathcal{S}_{1}$ in the recursion of Remark 3.1). To maintain the FCR control at level $\alpha=0.1$, InfoSP adjusts the width of the prediction interval in an accurate way to accommodate the informative constraint. InfoSCOP is roughly the same as InfoSP in this case, because the largest scores are kept in the calibration sample after initial selection. This is the most unfavorable setting for InfoSCOP, but it nevertheless performs similarly to InfoSP. In the second row, errors appear less often on the selection area, which makes the FCR of the naive selection and InfoSP too conservative. By contrast, the initial selection of InfoSCOP removes the largest scores of the calibration sample, resulting in much narrower prediction intervals, and thus in a much larger resolution-adjusted power (even better than the naive procedure).
@@
Figure 3: Informative prediction intervals when excluding $[a,b]$ (homoscedastic Gaussian regression model with perfect variance prediction), see text. First row: the predictor $\mu$ (dotted line) does not approximate well the true $\mu^{*}(x)=\mathbb{E}[Y|X=x]$ (solid line) in the selection area. Second row: the predictor $\mu$ (dotted line) does no approximate well the true $\mu^{*}$ (solid line) out of the selection area. The marginal prediction intervals are depicted in light blue, while the new informative prediction intervals (InfoSP and InfoSCOP) are depicted in red. While the plot corresponds to one data generation, the FCR and adjuster power computed in the title of each panel are computed with $100$ Monte-Carlo simulations. $n=1000$, $m=500$, $\alpha=0.1$.
@@
In Corollary 4.1, we consider the locally weighted score function $S_{y}(x)=|\mu(x)-y|/\sigma(x)$ for simplicity of exposition, but we can use any score function satisfying Assumption 1. For instance, for the quantile-based score function $S_{y}(x)=\max(q_{\beta_{0}}(x)-y,y-q_{\beta_{1}}(x))$, the corresponding $q_{i}$’s have the expression
$q_{i}=\overline{p}^{(a)}_{i}{\mathbf{1}\{\mu(X_{n+i})<a\}}+\overline{p}^{(b)}_{i}{\mathbf{1}\{\mu(X_{n+i})>b\}}+\overline{p}^{(\mu(X_{n+i}))}_{i}{\mathbf{1}\{a\leq\mu(X_{n+i})\leq b\}},$
where $\mu(x)=(q_{\beta_{0}}(x)+q_{\beta_{1}}(x))/2$ and where the $\overline{p}^{(y)}_{i}$’s are computed by using this score function. This leads to the prediction intervals $\mathcal{C}_{n+i}=[q_{\beta_{0}}(X_{n+i})-S_{(n_{\alpha}(\mathbf{p}))},q_{\beta_{1}}(X_{n+i})+S_{(n_{\alpha}(\mathbf{p}))}]$ (which do not contain $[a,b]$), by using the notation of Corollary 4.1.
@@ 4.2 Length-restricted prediction intervals
    InfoSP selects $\mathcal{S}=\mathrm{BH}(\mathbf{q})$ with $q_{i}$ given by (12).
@@
A similar result holds for InfoSCOP. In Corollary 4.2 (ii), the length of the prediction interval on the selection is always granted to be (at most) of the correct size, even if adjusting the level is necessary to account for selection (which de facto enlarge the prediction interval). Thanks to the $\mathrm{BH}(\mathbf{q})$ selection the size-adjustment is automatic, while maintaining the FCR control.
@@
By Remark 2.1, we have $q_{i}\leq\alpha$ iff $\sigma(X_{n+i})S_{(\lceil(1-\alpha)(n+1)\rceil)}\leq\lambda_{0}$. This implies the expression of $q_{i}$ and that $|\mathcal{C}_{n+i}|=2\sigma(X_{n+i})S_{(\lceil(1-\alpha)(n_{\alpha}(\mathbf{p})+1)\rceil)}\leq 2\lambda_{0}$ since $q_{i}\leq\alpha|\mathcal{S}(\mathbf{p})|/m$. ∎
@@
Corollary 4.2 easily extends to the case of conformalized quantile regression, by considering the quantile-based score function $S_{y}(x)=\max(q_{\beta_{0}}(x)-y,y-q_{\beta_{1}}(x))$. In that case, $q_{i}\leq\alpha$ iff $2S_{(\lceil(1-\alpha)(n+1)\rceil)}+q_{\beta_{1}}(X_{n+i})-q_{\beta_{0}}(X_{n+i})\leq 2\lambda_{0}$ which leads to the formula (13) for the $q_{i}$’s and to the prediction intervals $\mathcal{C}_{n+i}=[q_{\beta_{0}}(X_{n+i})-S_{(n_{\alpha}(\mathbf{p}))},q_{\beta_{1}}(X_{n+i})+S_{(n_{\alpha}(\mathbf{p}))}]$ (of length at most $2\lambda_{0}$), with the notation of Corollary 4.2.
@@
Figure 4: Informative prediction intervals when length-restricted (heteroscedastic Gaussian regression model with perfect mean prediction), see text. First row: the predictor $\sigma$ under-estimate the true $\sigma^{*}(x)=\mathbb{V}^{1/2}[Y|X=x]$ in the selection area. Second row: the predictor $\sigma$ over-estimate the true $\sigma^{*}$ in the selection area. The marginal prediction intervals are depicted in light blue, while the new informative prediction intervals (InfoSP and InfoSCOP) are depicted in red. While the plot corresponds to one data generation, the FCR and adjuster power computed in the title of each panel are computed with $100$ Monte-Carlo simulations. $n=1000$, $m=500$, $\alpha=0.1$.

### v3

@@ 3.2 Informative selective conditional prediction sets (InfoSCOP)
Throughout this section, we consider the iid model. It turns out that InfoSP can be too conservative in some contexts; this is manifest in the true FCR being much smaller than the nominal $\alpha$ level, see illustrations in § 4 and § 5. To avoid this, we can adapt the conditional approach of Bao et al., (2024) to our framework and start by selecting test samples and calibration samples. We would like, following the initial selection, to have as few as possible test samples for which $\mathcal{I}$-informative prediction sets cannot be constructed. For this purpose, we further split the calibration sample, which is a classical trick in conformal literature in order to preserve exchangeability with calibration samples after the initial selection (see Lemma H.5). Specifically, we follow the following steps
@@
    For $i+n\in\mathcal{S}^{(0)}\cap[n+1,n+m]$, compute the conformal $p$-values using calibration set $\{(X_{j},Y_{j}),j\in\mathcal{S}^{(0)}\cap[r+1,n]\}$ (i.e., using the conditional empirical distribution, post initial selection):
@@
We assume that the initial selection $\mathcal{S}^{(0)}$ satisfies the following permutation preserving assumption.
@@
The initial selection $\mathcal{S}^{(0)}$ is typically the result of a multiple testing procedure that is applied to the examples in $[r+1,n+m]$, that uses $p$-values computed with $((X_{j},Y_{j}),j\in[r])$ as calibration sample and $(X_{i},i\in[r+1,n+m])$ as covariate test sample, which immediately satisfies Assumption 3. We can use $\mathrm{BH}(\mathbf{q})$ as an initial selection stage (where the $q_{i}$’s are computed with the aforementioned sample split) so that selected examples are likely to correspond to $Y_{n+i}$ where an informative prediction set can be built. For excluding a null range in regression or a null class in classification, another choice is to use an appropriate BH procedure for testing that the examples from $[r+1,n+m]$ are from that null, see examples in § 4.1 and § 5.2.
@@
The informative selective conditional prediction set procedure pre-processed with the initial selection rule $\mathcal{S}^{(0)}$, called InfoSCOP, is defined as the InfoSP procedure of Definition 1 applied with the pre-processed $p$-value family $\mathbf{\overline{p}}^{0}$ (22), that is, $\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSCOP}}}(\mathbf{\overline{p}})=\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{\overline{p}}^{0})=(\mathcal{C}^{\alpha^{0}}_{n+i}(\mathbf{\overline{p}}^{0}))_{i\in\mathrm{BH}(\mathbf{q}^{0})},$ where $\alpha^{0}=\alpha|\mathrm{BH}(\mathbf{q}^{0})|/|\mathcal{S}^{(0)}\cap[n+1,n+m]|$ and the $\mathcal{I}$-adjusted $p$-values $\mathbf{q}^{0}$ are computed via (11) from the pre-processed $p$-values $\mathbf{\overline{p}}^{0}$.
@@
The proof, provided in § E.2, follows directly from the general calibration splitting trick Lemma H.5.
As we will see in the next sections, while it maintains the FCR guarantee, InfoSCOP can greatly improve over InfoSP. The main reason is that adjusting for selection is cheaper after the initial selection step: by reducing the fraction of examples in the test sample for which it is not possible to construct $\mathcal{I}$-informative prediction sets, the correction term $\alpha^{0}$ is expected to be close to $\alpha$ (or not much smaller than $\alpha$). Another reason is that the selection-conditional $p$-values $\mathbf{\overline{p}}^{0}$ will be better in settings where the initial selection step tends to remove the examples from the calibration set that have large non-conformity scores, see § 4, § I, and § J for such cases.
In general, the way InfoSCOP can improve over InfoSP depends on the context. For instance, for excluding a null range in classification, a null class in regression or for length restriction in regression, we show in § 4 and § 5, respectively, the great potential advantage of using the initial selection. On the other hand, we also show that for non-trivial classification, there may be no advantage of initial selection (in fact, there can be a slight disadvantage since the calibration sample after initial selection is smaller, as demonstrated in § 5.3).
@@ 4 Application to regression
This section is devoted to the regression case (that is, $\mathcal{Y}=\mathbb{R}$), as already introduced in Example 1.1. Throughout the section, we work in the iid model. Illustrations for other informative selections and an application for gene expression prediction from promoter sequences can be found in § I in the SM.
4.1 Excluding $[a,b]$ from the prediction interval
We focus here on the case where the user wants to build prediction intervals only for outcomes such that $Y_{n+i}<a$ or $Y_{n+i}>b$, which corresponds to excluding $\mathcal{Y}_{0}=[a,b]$ from the prediction interval, where $a<b$ are two benchmark values. This corresponds to common practice where users are interested only in reporting prediction intervals for individuals with “abnormal” outcomes. Setting $a=-\infty$ recovers the case where we only want to report prediction intervals for examples such that $Y_{n+i}>b$, which is the selection considered in Jin and Candes, (2023). We focus on two-sided prediction intervals here (the case of one-sided prediction intervals is postponed to § I). The choice of score function defines InfoSP and entails all the desired inferential guarantees. We formalize this for the locally weighted score function in Corollary 4.1. Using other score functions is also possible, e.g., the score function that corresponds to conformal quantile regression, see Remark 4.1.
@@ 4.1 Excluding $[a,b]$ from the prediction interval
Consider the iid model in the regression case, the locally weighted score function $S_{y}(x)=|\mu(x)-y|/\sigma(x)$ and suppose that Assumption 1 holds. Then the following holds for InfoSP with informative collection =$\mathcal{I}=\{\mbox{$I$ interval of }\mathbb{R}\>:\>I\cap[a,b]=\emptyset\}$ and full-calibrated $p$-value collection $\mathbf{\overline{p}}$ (7):
@@
    $q_{i}=\overline{p}^{(a)}_{i}{\mathbf{1}\{\mu(X_{n+i})<a\}}+\overline{p}^{(b)}_{i}{\mathbf{1}\{\mu(X_{n+i})>b\}}+{\mathbf{1}\{a\leq\mu(X_{n+i})\leq b\}}.$ (23)
@@
    The selected prediction intervals do not intersect $[a,b]$ and are of the form $\mathcal{C}_{n+i}=[\mu(x)-S_{(n_{\alpha}(\mathbf{p}))}\sigma(x),\mu(x)+S_{(n_{\alpha}(\mathbf{p}))}\sigma(x)]$, where $S_{(1)}\leq\dots\leq S_{(n)}$ are the ordered calibration scores $S_{Y_{j}}(X_{j})$, $1\leq j\leq n$ (with $S_{(n+1)}=+\infty$), and $n_{\alpha}(\mathbf{p})=\lceil(1-\alpha|\mathcal{S}(\mathbf{p})|/m)(n+1)\rceil$.
@@
Point (i) follows from straightforward computations, see Section F. Point (ii) follows from Lemma 2.1, (iii) from Remark 2.1 and the fact that $q_{i}\leq\alpha|\mathcal{S}(\mathbf{p})|/m$ iff $\mathcal{C}_{n+i}$ does not intersect $[a,b]$. Point (iv) follows from Theorem 3.1. ∎
Hence, our method, in addition to providing an FCR control on the selected ensures that the obtained prediction intervals are informative in the sense that they do not include benchmark values (i.e., values in $[a,b]$). This ensures that the selection method is meaningful for the considered aim, which formally entails the FDR control (ii). Obviously, since InfoSCOP is an InfoSP method for preprocessed $p$-values (22), a similar result holds for InfoSCOP, for any initial selection step $\mathcal{S}^{(0)}\subset[r+1,n+m]$ that satisfies the permutation preserving Assumption 3.
Corollary 4.1 is illustrated on Figure 2 when $\mathcal{S}^{(0)}$ is taken here as $\mathrm{BH}(\mathbf{q})$ at level $2\alpha$ (with the score $S_{y}(x)=|\mu(x)-y|/\sigma(x)$). In the first row, errors are larger further away from $[a,b]$. Hence, while the marginal prediction intervals control the FCR at level $\alpha$ when selecting all the covariates (as granted by (9)), the FCR is inflated for a naive selection that selects each example with a prediction interval at level $\alpha$ not intersecting $[a,b]$ (that is, naive selection is given by $\mathcal{S}_{1}$ in the recursion of Remark 3.1). To maintain the FCR control at level $\alpha=0.1$, InfoSP adjusts the width of the prediction interval in an accurate way to accommodate the informative constraint. InfoSCOP is roughly the same as InfoSP in this case, because the largest scores are kept in the calibration sample after initial selection. This is the most unfavorable setting for InfoSCOP, but it nevertheless performs similarly to InfoSP. In the second row, errors are smaller further away from $[a,b]$, which makes the FCR of the naive selection and InfoSP far too conservative. By contrast, the initial selection of InfoSCOP removes the largest scores of the calibration sample, resulting in much narrower prediction intervals, and thus in a much larger resolution-adjusted power (even better than the naive procedure).
@@
Figure 2: Informative prediction intervals when excluding $[a,b]$ (homoscedastic Gaussian regression model with perfect variance prediction), see text. The predictor $\mu$ (dotted line) does not approximate well the true $\mu^{*}(x)=\mathbb{E}[Y|X=x]$ (solid line) in the selection area (top row) and out of the selection area (bottom row). The marginal and informative prediction intervals (InfoSP and InfoSCOP) are depicted in light blue and red, respectively. While the plot corresponds to one data generation, the FCR and adjusted power computed in the title of each panel are computed with $100$ Monte-Carlo simulations. $n=1000$, $m=500$, $\alpha=0.1$.
@@
In Corollary 4.1, we consider the locally weighted score function $S_{y}(x)=|\mu(x)-y|/\sigma(x)$ for simplicity of exposition, but we can use any score function satisfying Assumption 1. For instance, for the quantile-based score function $S_{y}(x)=\max(q_{\beta_{0}}(x)-y,y-q_{\beta_{1}}(x))$, the corresponding $q_{i}$ have the expression $q_{i}=\overline{p}^{(a)}_{i}{\mathbf{1}\{\mu(X_{n+i})<a\}}+\overline{p}^{(b)}_{i}{\mathbf{1}\{\mu(X_{n+i})>b\}}+\overline{p}^{(\mu(X_{n+i}))}_{i}{\mathbf{1}\{a\leq\mu(X_{n+i})\leq b\}},$ where $\mu(x)=(q_{\beta_{0}}(x)+q_{\beta_{1}}(x))/2$ and where the $\overline{p}^{(y)}_{i}$’s are computed by using this score function (see the computations in Section F). This leads to the prediction intervals $\mathcal{C}_{n+i}=[q_{\beta_{0}}(X_{n+i})-S_{(n_{\alpha}(\mathbf{p}))},q_{\beta_{1}}(X_{n+i})+S_{(n_{\alpha}(\mathbf{p}))}]$ (which do not intersect $[a,b]$), by using the notation of Corollary 4.1.
@@ 4.2 Length-restricted prediction intervals
    InfoSP selects $\mathcal{S}=\mathrm{BH}(\mathbf{q})$ with $q_{i}$ given by the formula of Example 2.1.
@@
A similar result holds for InfoSCOP. In Corollary 4.2 (ii), the length of the prediction interval on the selection is always granted to be (at most) of the correct size $2\lambda_{0}$, even if adjusting the level is necessary to account for selection (which de facto enlarge the prediction interval). Thanks to the $\mathrm{BH}(\mathbf{q})$ selection the size-adjustment is automatic, while maintaining the FCR control.
@@
The expression of $q_{i}$ follows from straightforward computations, see Section F. This implies that $|\mathcal{C}_{n+i}|=2\sigma(X_{n+i})S_{(\lceil(1-\alpha)(n_{\alpha}(\mathbf{p})+1)\rceil)}\leq 2\lambda_{0}$ since $q_{i}\leq\alpha|\mathcal{S}(\mathbf{p})|/m$.
@@
Corollary 4.2 easily extends to the case of conformalized quantile regression, by considering the quantile-based score function $S_{y}(x)=\max(q_{\beta_{0}}(x)-y,y-q_{\beta_{1}}(x))$. In that case, the formula of the $q_{i}$’s are given in Example 2.1 (see Section F for a proof) and the prediction intervals are $\mathcal{C}_{n+i}=[q_{\beta_{0}}(X_{n+i})-S_{(n_{\alpha}(\mathbf{p}))},q_{\beta_{1}}(X_{n+i})+S_{(n_{\alpha}(\mathbf{p}))}]$ (of length at most $2\lambda_{0}$), with the notation of Corollary 4.2.
@@
Figure 3: Informative prediction intervals when length-restricted (heteroscedastic Gaussian regression model with perfect mean prediction), see text. The predictor $\sigma$ under-estimates (top row) and over-estimates (bottom row) the true $\sigma^{*}(x)=\mathbb{V}^{1/2}[Y|X=x]$ in the selection area. The marginal and informative prediction intervals (InfoSP and InfoSCOP) are depicted in light-blue and red, respectively. While the plot corresponds to one data generation, the FCR and adjusted power computed in the title of each panel are computed with $100$ Monte-Carlo simulations. $n=1000$, $m=500$, $\alpha=0.1$.